If a detonator is exploded on a railway line, an observer standing on the rail 3km away heard 2 sounds, evaluate the time interval between these sounds, the ypong modulus of steel is given as 2x1011N/m2, the density of steel 800kg/m3, the ratio of the heat capacity of air is given as 1.40


1
Expert's answer
2021-08-02T08:40:38-0400
"\Delta t=\left(\sqrt{\frac{\rho_a}{\gamma P}}-\sqrt{\frac{\rho_s}{Y_s}}\right)s\\
\Delta t=\left(\sqrt{\frac{1.4}{1.4\cdot10^5}}-\sqrt{\frac{8000}{2\cdot10^{11}}}\right)3000=8.89\ s"